OpenStudy (anonymous):

Convert x = sqrt(5)y to polar form

OpenStudy (anonymous):

x = r cos(theta) y = r sin(theta) r cos(theta) = sqrt(5) r sin(theta) cos(theta)/sin(theta) = sqrt(5) cot(theta) = sqrt(5)